log4x+log4(x+12)=3
(x)(x+12)=4^3
x^2+4x+12=48
x^2+4x=36
x(x+12)=36
x=?
I'm missing something right?

Answer by stanbon(75887)

log4x+log4(x+12)=3
(x)(x+12)=4^3
x^2+12x-64=0
(x+16)(x-4)=0
x=4
Cheers,
Stan H.